The mean of 25 observations is N,the total of the observations is 300,while the total sum is 625,find N,and the standard deviation of the observation
1
Expert's answer
2020-02-25T10:55:03-0500

The sample mean


"N={\displaystyle\sum_{i=1}^nx_i \over n}"

Given that "\displaystyle\sum_{i=1}^nx_i=300, n=25." Then


"N={300\over 25}=12"

The standard deviation of the observation


"\sigma=\sqrt{{\displaystyle\sum_{i=1}^n(x_i-N)^2 \over n}}"

Given that "\displaystyle\sum_{i=1}^n(x_i-N)^2=625." Then


"\sigma=\sqrt{{625 \over 25}}=5"

The mean of 25 observations is "N=12."

The standard deviation of the observation is "\sigma=5."